# Dual topological insulator with mirror-symmetry-protected helical edge states

### v1

@@ VI Conclusions
We have obtained the electronic band structure of the quasi-2D ternary Na2CdSn tri-atomic layer via DFT calculations and derived an effective $4\times 4$ model using the $\bm{k}\cdot\bm{p}$ method to describe the low-energy states. To do so, we have employed the Lödwin perturbation theory, the folding down technique, and the theory of invariants. Our rigorously derived effective Hamiltonian contrasts with that of Ref. [70]. In particular, it has some quadratic off-diagonal terms (absent in [70]) that are crucial to obtain our spin and mirror Chern numbers and the corresponding (two pairs of) protected edge states.
More specifically, by fitting our effective model to the DFT energy bands, we have obtained numerical values for the model parameters and predicted that Na2CdSn is a dual topological insulator (DTI) with spin and mirror Chern numbers given by $C_{s}=-2$ and $C_{m}=-2$, respectively. Using the envelope function approach, we have also obtained analytical dispersion relations, quadratic in the crystal momentum, for the topological edge states in the semi-infinite plane geometry. The edge modes have also been analyzed numerically in the ribbon geometry. For both geometries, we have found two pairs of counter-propagating topological edge states at each boundary, as expected from the bulk-boundary correspondence. Both our Chern numbers and the corresponding (number of) edges states are in blatant contrast to the results in Ref. [70]; there it was reported spin and mirror Chern numbers $-1$, and consequently, only one pair of topological edge states.
Our investigation contributes to the understanding of Na2CdSn as a giant gap ($E_{g}=234.8\,$meV) DTI in 2D by providing its proper model Hamiltonian and the corresponding effective $\bm{k}\cdot\bm{p}$ parameters. This enables us to derive the energy dispersions and wave functions of the edge states arising in finite geometries of this material. The Na2XY (X=Mg,Cd;Y=Pb,Sn) family of tri-atomic layers are the building stacking blocks of 3D van der Waals Dirac semimetals, already synthesized experimentally [46, 45]. They can also be potentially obtained via mechanical exfoliation. The DTI Na2CdSn represents a promising alternative to conventional small-gap topological insulators in 2D – which usually require the engineering of elaborated quantum wells [5, 6, 99, 108, 58]. We are optimistic that our findings will stimulate the experimental verification of the DTI phase in this outstanding candidate for realistic room-temperature applications in nanotechnology, such as nanoelectronic, spintronic, thermoelectric and optical devices.
@@ VII Acknowledgments
W. H. C. acknowledges helpful discussions with Michael Flatté and Jaroslav Fabian. This work was supported by the National Council for Scientific and Technological Development (CNPq) Grant No. 301595/2022-4, PDJ Process Number 152321/2020-9 and Grant No. 306122/2018-9; by Coordenação de Aperfeiçoamento de Pessoal de Nível Superior (PNPD/CAPES); by São Paulo Research Foundation (FAPESP), Grant 2020/00841-9. J.Z. acknowledges the financial support of the Quantimony (Marie Sklodowska-Curie project Grant 956548). P.E.F.J and W.H.C. acknowledge the financial support of the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) SFB 1277 (Project-ID 314695032, projects B07 and B11). P.E.F.J. acknowledges the financial support of the DFG-SPP 2244 (Project No. 443416183).
Appendix A The mirror symmetry operation
To properly define the mirror symmetry operation about the $z$-axis (in this case the mirror plane is the $xy$-plane), we consider its effect on both spin and spatial degrees of freedom, since spinful electrons live in the state space given by the direct product between the spin and orbital subspaces [109, 110]. The spin angular momentum transforms as an axial vector, and therefore, the components of the spin operator perpendicular and parallel to the mirror plane must satisfy, respectively,
@@ Appendix A The mirror symmetry operation
with $i=x,y$ [111]. The mirror symmetry operator, consistent with (52), is given by the direct product
@@
The $\hat{\sigma}_{h}^{spin}$ operator can be defined as a $\pi$-rotation about the $z$ axis followed by the inversion operation (both in the spin subspace) [32, 20, 112], that is,
@@
$\hat{R}_{z}^{spin}(\pi)=e^{-i\frac{\hat{S}_{z}}{\hbar}\pi}=-i\hat{\sigma}_{z}.$ (55)
Hence the mirror operator in the state space [Eq. (53)] reads
@@ Appendix B Expressions for the model parameters in the Löwdin perturbation approach
In Subsec. III.1 we have shown the effective $\bm{k}\cdot\bm{p}$ Hamiltonian derived using the Löwdin perturbation theory. Below, we show the explicit expressions for the parameters of the model. The parameters $A_{1}$ and $A_{2}$ are given, respectively, by
$A_{1}=\left(\frac{\hbar}{m_{0}}\right)^{2}\left[\sum_{\beta}^{B[\Gamma_{1}]}\frac{|\langle X|p_{x}|\Gamma_{1\beta}\rangle|^{2}}{\epsilon_{0}-E_{\Gamma_{1\beta}}}+\sum_{\beta}^{B[\Gamma_{6}]}\frac{|\langle X|p_{x}|\Gamma_{6\beta}^{x}\rangle|^{2}}{\epsilon_{0}-E_{\Gamma_{6\beta}}}\right]\qquad\text{and}$ (60)
@@
where we have used $\langle X|p_{y}|\Gamma_{6\beta}^{y}\rangle=-\langle X|p_{x}|\Gamma_{6\beta}^{x}\rangle$ (similar relations were used to simplify all the matrix elements below). In the expressions above, $B[\Gamma_{i}]$ indicates that $\beta$ runs over all infinite bands belonging to the $\Gamma_{i}$ IRREP. When the IRREP of the $\beta$-th band is one-dimensional, its single basis function is written as $|\Gamma_{i\beta}\rangle$. For IRREPs of higher dimensions, we use $|\Gamma_{i\beta}^{\mu}\rangle$ to denote the basis function that transforms according to $\mu$, i.e. $|\Gamma_{i\beta}^{\mu}\rangle\sim\mu$, $(\mu=x,y,R_{x},R_{y}...)$. $E_{\Gamma_{i\beta}}$ is the energy of the $\beta$-th band belonging to the $\Gamma_{i}$ IRREP. Naturally, all basis functions of an IRREP have the same energy eigenvalue, and therefore, it is not necessary to label the energy eigenvalue with the basis function. The remaining parameters read
@@ Appendix C Model derivation via the theory of invariants
and finding the unknown coefficients $h_{l,m,n}$ (matrices) by imposing that the Hamiltonian $\mathcal{H}$ of the crystal, characterized by a point group $G$, be invariant under all symmetry operations $g_{i}$ ($g_{i}\in G$), i.e., $[\mathcal{H},g_{i}]=0$. For the Bloch Hamiltonian $\mathcal{H}(\bm{k})=e^{-i\bm{k}\cdot\bm{r}}\mathcal{H}e^{i\bm{k}\cdot\bm{r}}$, this commutator yields
@@
where $u_{g_{i}}$ is the matrix representation of $g_{i}$ in $\bm{k}$ space. An equivalent expression can be derived for the projected Hamiltonian $H(\bm{k})$ in (69), given by
@@
with $U_{g_{i}}$ the matrix representation of $g_{i}$ in the Hilbert space. Hence, by using (69) and the constraints imposed by (71), we obtain a system of equations for the coefficients $h_{l,m,n}$. Note that for 2D systems, which is the case we are interested in here, we can simply drop $k_{z}$ and the sum over $n$ in Eq. (69).
The Na2CdSn tri-atomic layer we are interested in belongs to the point group $D_{3h}$. To determine $h_{l,m,n}$, it is sufficient to use in (71) only the corresponding group generators, $\{g_{i}\}=\{C_{3z},C_{2x},\sigma_{h}\}$, with $C_{3z}$ representing a rotation about the $z$-axis, $C_{2x}$ a rotation about the $x$-axis, and $\sigma_{h}$ a mirror plane reflection about the $xy$-plane (see character table in the SM [76]). In addition, we take into account that the system is time-reversal-symmetric.
Here we use the QSymm Python package [113], which receives as input the matrix representations $U_{g_{i}}$ and ${g_{i}}$, for all $g_{i}\in G$ (generators + TRS), and returns the matrices $h_{l,m,n}$. The matrix representations $U_{g_{i}}$ of the generators $\{C_{3z},C_{2x},\sigma_{h}\}$ and time reversal operator $\mathcal{T}$ follow from both the two-dimensional IRREP of the orbital part ($\Gamma_{6}$) and the “spinorial” part ($\Gamma_{7}$), i.e., $\Gamma_{6}\otimes\Gamma_{7}=\Gamma_{8}\oplus\Gamma_{9}$, and are given by
$U_{C_{3z}}$ $=$ $e^{-i\phi\sigma_{3}/2}\otimes e^{-i\phi\tau_{2}},$ (72)
@@
with $\phi=2\pi/3$. Up to second-order in the crystal momentum $\bm{k}$, the $4\times 4$ Hamiltonian is given by
@@
$+$ $\begin{pmatrix}c_{4}k_{y}^{2}+c_{5}k_{x}^{2}&+(c_{5}-c_{4})k_{x}k_{y}+ic_{6}k^{2}&0&0\\ (c_{5}-c_{4})k_{x}k_{y}-ic_{6}k^{2}&c_{4}k_{x}^{2}+c_{5}k_{x}^{2}&0&0\\ 0&0&c_{4}k_{y}^{2}+c_{5}k_{x}^{2}&(c_{5}-c_{4})k_{x}k_{y}-ic_{6}k^{2}\\ 0&0&(c_{5}-c_{4})k_{x}k_{y}+ic_{6}k^{2}&c_{4}k_{x}^{2}+c_{5}k_{y}^{2}\end{pmatrix},$

### v2

@@ VI Conclusions
We have obtained the electronic band structure of the quasi-2D ternary Na2CdSn triatomic layer via DFT calculations and derived an effective $4\times 4$ model using the $\bm{k}\cdot\bm{p}$ method to describe the low-energy states. To do so, we have employed the Lödwin perturbation theory, the folding-down technique, and the theory of invariants. Our rigorously derived effective Hamiltonian contrasts with that of Ref. [70]. It has some quadratic off-diagonal terms (absent in Ref. [70]) that are crucial to obtain our spin and mirror Chern numbers and the corresponding (two pairs of) protected edge states.
More specifically, by fitting our effective model to the DFT energy bands, we have obtained numerical values for the model parameters and predicted that Na2CdSn is a DTI with spin and mirror Chern numbers given by $C_{s}=-2$ and $C_{m}=-2$, respectively. Using the envelope function approach, we have also obtained analytical dispersion relations, quadratic in the crystal momentum, for the topological edge states in the semi-infinite plane geometry. The edge modes have also been analyzed numerically in the ribbon geometry. For both geometries, we have found two pairs of counterpropagating topological edge states at each boundary, as expected from the bulk-boundary correspondence. Both our Chern numbers and the corresponding (number of) edge states are in blatant contrast with the results in Ref. [70]; there, spin and mirror Chern numbers $-1$ were reported and, consequently, only one pair of topological edge states.
[…]
@@ VII Acknowledgments
W.H.C. acknowledges helpful discussions with Michael Flatté and Jaroslav Fabian. This paper was supported by the National Council for Scientific and Technological Development (CNPq) Grant No. 301595/2022-4, PDJ Process No. 152321/2020-9, and Grant No. 306122/2018-9; by Coordenação de Aperfeiçoamento de Pessoal de Nível Superior (PNPD/CAPES); by São Paulo Research Foundation (FAPESP), Grant No. 2020/00841-9. J.Z. acknowledges the financial support of the Quantimony (Marie Sklodowska-Curie Project Grant No. 956548). P.E.F.J and W.H.C. acknowledge the financial support of the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) SFB 1277 (Project-ID No. 314695032, Projects No. B07 and No. B11). P.E.F.J. acknowledges the financial support of the DFG-SPP 2244 (Project No. 443416183). W.H.C. acknowledges funding from the Deutsche Forschungsgemeinschaft (DFG) Grant No. TRR 173 268565370 (Project A03).
Appendix A Mirror symmetry operation
To properly define the mirror symmetry operation about the $z$ axis (in this case, the mirror plane is the $xy$ plane), we consider its effect on both spin and spatial degrees of freedom since spinful electrons live in the state space given by the direct product between the spin and orbital subspaces [113, 114]. The spin angular momentum transforms as an axial vector, and therefore, the components of the spin operator perpendicular and parallel to the mirror plane must satisfy, respectively,
@@ Appendix A Mirror symmetry operation
with $i=x,y$ [115]. The mirror symmetry operator, consistent with Eq. (52), is given by the direct product:
@@
The $\hat{\sigma}_{h}^{spin}$ operator can be defined as a $\pi$ rotation about the $z$ axis followed by the inversion operation (both in the spin subspace) [32, 20, 116], that is,
@@
$\hat{R}_{z}^{spin}(\pi)=\exp\left(-i\frac{\hat{S}_{z}}{\hbar}\pi\right)=-i\hat{\sigma}_{z},$ (55)
where we have used $\hat{S}_{z}=\frac{\hbar}{2}\hat{\sigma}_{z}$. Hence, the mirror operator in the state space [Eq. (53)] reads
@@ Appendix B Expressions for the model parameters in the Löwdin perturbation approach
In Sec. III.1, we have shown the effective $\bm{k}\cdot\bm{p}$ Hamiltonian derived using the Löwdin perturbation theory. Below, we show the explicit expressions for the parameters of the model. The parameters $A_{1}$ and $A_{2}$ are given, respectively, by
$A_{1}=\left(\frac{\hbar}{m_{0}}\right)^{2}\left[\sum_{\beta}^{B[\Gamma_{1}]}\frac{|\langle X|p_{x}|\Gamma_{1\beta}\rangle|^{2}}{\epsilon_{0}-E_{\Gamma_{1\beta}}}+\sum_{\beta}^{B[\Gamma_{6}]}\frac{|\langle X|p_{x}|\Gamma_{6\beta}^{x}\rangle|^{2}}{\epsilon_{0}-E_{\Gamma_{6\beta}}}\right],$ (60)
and
@@
where we have used $\langle X|p_{y}|\Gamma_{6\beta}^{y}\rangle=-\langle X|p_{x}|\Gamma_{6\beta}^{x}\rangle$ (similar relations were used to simplify all the matrix elements below). In the expressions above, $B[\Gamma_{i}]$ indicates that $\beta$ runs over all infinite remote bands belonging to the $\Gamma_{i}$ irrep. When the irrep of the $\beta$th band is 1D, its single basis function is written as $|\Gamma_{i\beta}\rangle$. For irreps of higher dimensions, we use $|\Gamma_{i\beta}^{\mu}\rangle$ to denote the basis function that transforms according to $\mu$, i.e., $|\Gamma_{i\beta}^{\mu}\rangle\sim\mu$, $(\mu=x,y,R_{x},R_{y}...)$. Here, $E_{\Gamma_{i\beta}}$ is the energy of the $\beta$th band belonging to the $\Gamma_{i}$ irrep. Naturally, all basis functions of an irrep have the same energy eigenvalue, and therefore, it is not necessary to label the energy eigenvalue with the basis function. The remaining parameters read
@@ Appendix C Model derivation via the Theory of Invariants
and finding the unknown coefficients $h_{l,m,n}$ (matrices) by imposing that the Hamiltonian $\mathcal{H}$ of the crystal, characterized by a point group $G$, be invariant under all symmetry operations $g_{i}$ ($g_{i}\in G$), i.e., $[\mathcal{H},g_{i}]=0$. For the Bloch Hamiltonian $\mathcal{H}(\bm{k})=\exp(-i\bm{k}\cdot\bm{r})\mathcal{H}\exp(i\bm{k}\cdot\bm{r})$, this commutator yields
@@
where $u_{g_{i}}$ is the matrix representation of $g_{i}$ in $\bm{k}$ space. An equivalent expression can be derived for the projected Hamiltonian $H(\bm{k})$ in Eq. (69), given by
@@
with $U_{g_{i}}$ the matrix representation of $g_{i}$ in the Hilbert space. Hence, by using Eq. (69) and the constraints imposed by Eq. (71), we obtain a system of equations for the coefficients $h_{l,m,n}$. Note that, for 2D systems, which is the case we are interested in here, we can simply drop $k_{z}$ and the sum over $n$ in Eq. (69).
The Na2CdSn triatomic layer we are interested in belongs to the point group $D_{3h}$. To determine $h_{l,m,n}$, it is sufficient to use in Eq. (71) only the corresponding group generators $\{g_{i}\}=\{C_{3z},C_{2x},\sigma_{h}\}$, with $C_{3z}$ representing a rotation about the $z$ axis, $C_{2x}$ a rotation about the $x$ axis, and $\sigma_{h}$ a mirror plane reflection about the $xy$ plane (see character table in the SM [75]). In addition, we consider that the system is time-reversal symmetric.
Here, we use the QSYMM Python package [117], which receives as input the matrix representations $U_{g_{i}}$ and ${g_{i}}$, for all $g_{i}\in G$ (generators + TRS), and returns the matrices $h_{l,m,n}$. The matrix representations $U_{g_{i}}$ of the generators $\{C_{3z},C_{2x},\sigma_{h}\}$ and time-reversal operator $\mathcal{T}$ follow from both the 2D irrep of the orbital part ($\Gamma_{6}$) and the spinorial part ($\Gamma_{7}$), i.e., $\Gamma_{6}\otimes\Gamma_{7}=\Gamma_{8}\oplus\Gamma_{9}$, and are given by
$U_{C_{3z}}$ $=$ $\exp\left(-\frac{i\phi\sigma_{3}}{2}\right)\otimes\exp(-i\phi\tau_{2}),$ (72)
@@
with $\phi=2\pi/3$. Up to second order in the crystal momentum $\bm{k}$, the $4\times 4$ Hamiltonian is given by
@@
$+$ $\begin{pmatrix}c_{4}k_{y}^{2}+c_{5}k_{x}^{2}&(c_{5}-c_{4})k_{x}k_{y}+ic_{6}k^{2}&0&0\\ (c_{5}-c_{4})k_{x}k_{y}-ic_{6}k^{2}&c_{4}k_{x}^{2}+c_{5}k_{x}^{2}&0&0\\ 0&0&c_{4}k_{y}^{2}+c_{5}k_{x}^{2}&(c_{5}-c_{4})k_{x}k_{y}-ic_{6}k^{2}\\ 0&0&(c_{5}-c_{4})k_{x}k_{y}+ic_{6}k^{2}&c_{4}k_{x}^{2}+c_{5}k_{y}^{2}\end{pmatrix},$